OpenStudy (anonymous):

4(2x-3) = 3(5x + 1) + 2

OpenStudy (curry):

8x-12=15x+5

OpenStudy (curry):

-7x=17

OpenStudy (curry):

x=-17/7

OpenStudy (anonymous):

8x - 12 = 15x + 5 x = -17/7

OpenStudy (anonymous):

Thanks a lot :)

OpenStudy (anonymous):

8x-12=15x+3+2 -5-12=15x-8x 7x=-17 x=-17/7

OpenStudy (curry):

dunnie instead of thanking show your appreciation with a medal